OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (anonymous):

okayokayokayokayokay I'mI'mI'mI'm notnotnotnot thatthatthatthatthat sursuresuresuresure

OpenStudy (anonymous):

Here is the point slope form equation of a line: y - y1 = m(x - x1) (x1, y1) is a known point, m is the slope of the line, and(x, y) is any other point on the line. Plug in your known slope and point, (-1,6) into the point slope form equation of a line. This is what i got: y - 6 = -3(x + 1) Let me know what you get, I could be wrong!

OpenStudy (anonymous):

yeahyeahyeah whatwhatwhat hehehe ororor shesheshe saidsaidsaid

OpenStudy (anonymous):

why did it change to addition at the end of the equation , (x+1) ?

OpenStudy (anonymous):

if i followed slope point form it would be A so i'd like to know why it changed to subtraction @Swaim

OpenStudy (anonymous):

i ment addition